### v4

# Orthogonality questions in the Hardy space related to $\zeta$-zeros

Abstract.
A Hardy space approach to the Nyman-Beurling and Báez-Duarte criterion for the Riemann Hypothesis (RH) was introduced recently in [19] and further developed in [14]. It states that the RH holds if and only if a particular sequence of functions $(h_{k})_{k\geq 2}$ is complete in the Hardy space $H^{2}$. This article is concerned with orthogonality questions related to the family $(h_{k})_{k\geq 2}$. The first goal is to analyze the orthogonal complement of $\mathcal{N}=\mathrm{span}(h_{k})_{k\geq 2}$ in $H^{2}$. Unbounded Toeplitz operators on $H^{p}$ spaces and de Branges-Rovnyak spaces play a central role and our results show that the size and dimension of $\mathcal{N}^{\perp}$ reveal information on the zeros of the Riemann $\zeta$-function. The second goal is to show that $(h_{k})_{k\geq 2}$ possesses a complete biorthogonal sequence in $H^{2}$. We also discuss a folklore conjecture about the number of $\zeta$-zeros if the RH fails.
Key words and phrases:
Riemann hypothesis, Hardy space, local Dirichlet space, de Branges-Rovnyak space, Smirnov class.
2010 Mathematics Subject Classification:
Primary; Secondary
Introduction
A classical result of Nyman and Beurling (see [6], [20]) shows that the Riemann Hypothesis (RH) is equivalent to the completeness of $\{\rho_{\lambda}:0\leq\lambda\leq 1\}$ in $L^{2}(0,1)$ where $\rho_{\lambda}(x)=\{\lambda/x\}-\lambda\{1/x\}$ and $\{x\}$ denotes the fractional part of $x$. This is furthermore equivalent to the characteristic function $\chi_{(0,1)}$ belonging to the closed linear span of $\{\rho_{\lambda}:0\leq\lambda\leq 1\}$. Half a century later Báez-Duarte [2] strenghthened this result by showing that RH is true if and only if $\chi_{(0,1)}\in\overline{\mathrm{span}\{\rho_{1/k}:k\geq 2\}}$. See the expository article of Bagchi [3] and the survey of Balazard [4]. Recently the Nyman-Beurling and Báez-Duarte approaches to the RH have been explored via tools from Hardy space $H^{2}$ theory [19] and other analytic function spaces [14].
For each $k\geq 2$, define
$h_{k}(z)=\dfrac{1}{1-z}\log\left(\dfrac{1+z+\dots+z^{k-1}}{k}\right)$
and denote by $\mathcal{N}$ the linear span of $\{h_{k}:k\geq 2\}$. That each $h_{k}$ belongs to $H^{2}$ was proved in [19, Lemma 7]. One of the main results of [19] was a reformulation of Báez-Duarte’s result as a completeness problem in $H^{2}$. Then in [14] the same completeness problem in the $H^{p}$ spaces was shown to provide zero-free half planes for the Riemann $\zeta$-function. We state both results here as one.
Theorem 1.
The RH holds if and only if $\mathcal{N}$ is dense in $H^{2}$. If $\mathcal{N}$ is dense in $H^{p}$ for some $1<p\leq 2$, then
$\zeta(s)\neq 0\ \ \mathrm{for}\ \ \Re(s)>\frac{1}{p}.$
The density of $\mathcal{N}$ in $H^{1}$ gives the known zero-free region $\Re(s)\geq 1$.
Although $\mathcal{N}$ is unconditionally dense in $H^{p}$ for $10<p<$ 1 (see [14, Cor. 4.6]), this provides no new information regarding $\zeta$-zeros. Also note that the RH is equivalent to $\mathcal{N}^{\perp}=\{0\}$ in $H^{2}$ by Theorem 1. This suggests the question of the size and dimension of $\mathcal{N}^{\perp}$ and their possible relation to the $\zeta$-zeros. The first archetypal result addressing this is the following (See [19, Thm. 12]).
Theorem 2.
$\mathcal{N}^{\perp}\cap\mathcal{D}_{{\delta}_{1}}=\{0\},$ where $\mathcal{D}_{{\delta}_{1}}$ denotes the local Dirichlet space at $1$.
Since $\mathcal{D}_{{\delta}_{1}}\subset H^{2}\subset H^{p}$ for $20<p<$ 2, Theorems 1 and 2 inspire the following question: which linear spaces of analytic functions on $\mathbb{D}$ with $X_{1}\subset H^{2}\subset X_{2}$ satisfy
$(1)\ \mathcal{N}^{\perp}\cap X_{1}=\{0\}\ \ \ \mathrm{and}\ \ \ (2)\ \mathcal{N}\ \mathrm{is}\ \mathrm{dense}\ \mathrm{in}\ X_{2}\ ?$
Interestingly, both $(1)$ and $(2)$ are equivalent if one takes $X_{2}=X$ a Frechet space and $X_{1}=X^{*}$ its Cauchy dual (see Proposition 9). In Section 2 we analyze this orthogonality question using tools from de Branges-Rovnyak spaces and unbounded Toeplitz operators on $H^{p}$ (see Proposition 11), and in particular develop a new approach for finding zero-free half-planes for the $\zeta$ function (see Theorem 14).
In Section 3 we deal with the biorthogonality question: Does the sequence $(h_{k})_{k\geq 2}$ posses a complete biorthogonal sequence in $H^{2}$? (see Subsection $1.5$). We affirmatively answer this question (see Theorem 15). To provide some context, Vasyunin showed in [22] that the Báez-Duarte sequence $\{\rho_{1/k}:k\geq 2\}$ is minimal by constructing a biorthogonal sequence for it in $L^{2}(0,1)$. Whereas the sequence $\{\rho_{1/k}:k\geq 2\}$ is not complete in $L^{2}(0,1)$, the completeness of $(h_{k})_{k\geq 2}$ in $H^{2}$ is equivalent to the RH by Theorem 1. The property of completeness of a sequence is not in general inherited by its biorthogonal sequence. But it may do so in very special cases such as for sequences of complex exponentials in $L^{2}(-\pi,\pi)$ (see [25]). Therefore an affirmative answer to this question (independent of RH) is intriguing.
Finally in Section 4 we discuss a folklore conjecture about the $\zeta$-zeros which we name the RH failure (RHF) conjecture: If the RH fails, then it fails infinitely often. More precisely, either $\zeta$ has no nontrivial zeros outside the critical line, or it has infinitely many. The main result of this section (Theorem 21) states that
[…]
We were unable to locate a reference for this conjecture in the literature. But an informative discussion of this conjecture does appear in mathoverflow.net [11].
1. Preliminaries
We denote by $\mathbb{D}$ and $\mathbb{T}$ the open unit disk and the unit circle respectively. By $\mathrm{Hol}(\mathbb{D})$ we denote the space of holomorphic functions on $\mathbb{D}$ and the shift operator is defined by $(Sf)(z)=zf(z)$ for $f\in\mathrm{Hol}(\mathbb{D})$.
1.1. Hardy spaces and the Smirnov class
A holomorphic function $f$ on $\mathbb{D}$ belongs to the Hardy-Hilbert space $H^{2}$ if
[…]
The space $H^{2}$ is a Hilbert space with the $\ell^{2}$-inner product
$\langle f,g\rangle=\sum_{n=0}^{\infty}a_{n}\overline{b_{n}},$
where $(a_{n})_{n\in\mathbb{N}}$ and $(b_{n})_{n\in\mathbb{N}}$ are the Fourier coefficients for $f$ and $g$ respectively. For any $f\in H^{2}$ and $\zeta\in\mathbb{T}$, the radial limit $f^{*}(\zeta):=\lim_{r\to 1^{-}}f(r\zeta)$ exists $m$-a.e. on $\mathbb{T}$, where $m$ denotes the normalized Lebesgue measure on $\mathbb{T}$. Analogously for $p>0$, the Hardy space $H^{p}$ consists of those holomorphic $f$ on $\mathbb{D}$ such that
$\|f\|_{H^{p}}^{p}=\sup_{0<r<1}\int_{\mathbb{T}}\left|f(rz)\right|^{p}\,\mathrm{d}m(z)\,<\,\infty.$
The $H^{p}$ are Banach spaces for $p\geq 1$ and complete metric spaces for $10<p<$ 1 and $H^{\infty}$ denotes the space of bounded holomorphic functions on $\mathbb{D}$. A function $f$ is called a cyclic vector for the shift $S$ in $H^{p}$ if $\mathrm{span}(S^{n}f)_{n\geq 0}=\mathbb{C}[z]f$ is dense in $H^{p}$. When $p=2$ these cyclic vectors are commonly known as outer functions. Since $H^{\infty}\subset H^{p}$ for all $p>0$, the $H^{\infty}$ outer functions are cyclic for all $H^{p}$ spaces. The Smirnov class $N^{+}$ consists of all holomorphic functions $g/h$ on $\mathbb{D}$ such that $g,h\in H^{\infty}$ and $h$ is an outer function. The space $N^{+}$ is a topological algebra with respect to pointwise multiplication and $g/h\in N^{+}$ is a unit if both $g$ and $h$ are outer functions. The topology on $N^{+}$ can be metrized with the translation-invariant complete metric
$d(f,g)=\int_{\mathbb{T}}\log(1+\left|f-g\right|)\,\mathrm{d}m,\qquad f,g\in N^{+}.$
Similar to $H^{p}$ spaces, convergence in $N^{+}$ implies locally unifom convergence on $\mathbb{D}$ and functions have radial limits $m$-a.e. on $\mathbb{T}$. In fact we have $H^{p}\subset H^{q}\subset N^{+}$ for all $0<q<p\leq\infty$. Duren [10] is a classic reference for the $H^{p}$ and $N^{+}$ spaces.
1.2. Local Dirichlet spaces
Let $\mu$ be a finite positive Borel measure on $\mathbb{T}$, and let $P\mu$ denote its Poisson integral. The generalized Dirichlet space $\mathcal{D}_{\mu}$ consists of $f\in H^{2}$ satisfying
[…]
Then $\mathcal{D}_{\mu}$ is a Hilbert space with norm $\left\|f\right\|_{\mathcal{D}_{\mu}}^{2}:=\left\|f\right\|_{2}^{2}+\mathcal{D}_{\mu}(f)$. If $\mu=m$, then $\mathcal{D}_{m}$ is the classicial Dirichlet space. If $\mu=\delta_{\zeta}$ is the Dirac measure at $\zeta\in\mathbb{T}$, then $\mathcal{D}_{\zeta}:=\mathcal{D}_{\delta_{\zeta}}$ is called the local Dirichlet space at $\zeta$ and in particular
[…]
The recent book [18] contains a comprehensive treatment of local Dirichlet spaces and the following result establishes a criterion for their membership.
Theorem 3.
(See [18, Thm. 7.2.1]) Let $\zeta\in\mathbb{T}$ and $f\in\mathrm{Hol}(\mathbb{D})$. Then $\mathcal{D}_{\delta_{\zeta}}(f)<\infty$ if and only if
$f(z)=(z-\zeta)g(z)+a$
for some $g\in H^{2}$ and $a\in\mathbb{C}$. In particular $f^{*}(\zeta)$ exists for all $f\in\mathcal{D}_{\zeta}$.
So each local Dirichlet space $\mathcal{D}_{\zeta}=(S-\zeta I)H^{2}+\mathbb{C}$ is a proper subspace of $H^{2}$. We define the $H^{p}$-analogues of these spaces for $p>0$ by
$\mathcal{D}_{\zeta}^{p}:=(S-\zeta I)H^{p}+\mathbb{C}$
and note that $\mathcal{D}_{\zeta}^{2}=\mathcal{D}_{\delta_{\zeta}}$ and $\mathcal{D}_{\zeta}^{p}\subsetneq\mathcal{D}_{\zeta}^{q}$ for $q<p$ since $H^{p}\subsetneq H^{q}$. Straightforward but lengthy computations show that $\mathcal{D}_{\zeta}^{p}\subsetneq H^{2}$ for $p>1$ and $H^{2}\subsetneq\mathcal{D}_{\zeta}^{p}$ for $0<p<\frac{2}{3}$.
1.3. The de Branges-Rovnyak spaces
Given $\psi\in L^{\infty}(\mathbb{T})$, the corresponding Toeplitz operator $T_{\psi}:H^{2}\to H^{2}$ is defined by
$T_{\psi}f:=P_{+}(\psi f)$
where $P_{+}:L^{2}(\mathbb{T})\to H^{2}$ denotes the orthogonal projection of $L^{2}(\mathbb{T})$ onto $H^{2}$. Clearly $T_{\psi}$ is a bounded operator on $H^{2}$ with $||T_{\psi}||\leq||\psi||_{L^{\infty}}$. If $h\in H^{\infty}$, then $T_{h}$ is the operator of multiplication by $h$ and its adjoint is $T_{\overline{h}}$. Given $b$ in the closed unit ball of $H^{\infty}$, the de Branges-Rovnyak space $\mathcal{H}(b)$ is the image of $H^{2}$ under the operator $(I-T_{b}T_{\overline{b}})^{1/2}$. The general theory of $\mathcal{H}(b)$ spaces divides into two distinct cases, according to whether $b$ is an extreme point or a non-extreme point of the unit ball of $H^{\infty}$. We shall be concerned only with the non-extreme case. In this case there exists a unique outer function $a\in H^{\infty}$ such that $a(0)>0$ and $|a^{*}|^{2}+|b^{*}|^{2}=1$ a.e. on $\mathbb{T}$. The pair $(b,a)$ is called a Pythagorean pair and the function $b/a$ belongs to the Smirnov class $N^{+}$. That all $N^{+}$ functions arise as the quotient of a pair associated to a non-extreme function was shown by Sarason [21]. The two-volume work ([12][13]) is an encyclopedic reference for these spaces.
If $\varphi$ is a rational function in $N^{+}$ the corresponding pair $(b,a)$ is also rational (see [21, Remark. 3.2]). Constara and Ransford [8] characterized the rational pairs $(b,a)$ for which $\mathcal{H}(b)$ is a generalized Dirichlet space.
Theorem 4.
(See [8, Theorem 4.1]) Let $(b,a)$ be a rational pair and $\mu$ a finite positive measure on $\mathbb{T}$. Then $\mathcal{H}(b)=\mathcal{D}_{\mu}$ if and only if
  1. (1)
    the zeros of $a$ on $\mathbb{T}$ are all simple, and
  2. (2)
    the support of $\mu$ is exactly equal to this set of zeros.
As an example, if $(b,a)$ is the rational pair associated with the $N^{+}$ function $\varphi(z)=\frac{1}{1-\zeta}$ for $\zeta\in\mathbb{T}$, then $\mathcal{H}(b)=\mathcal{D}_{\zeta}$ is a local Dirichlet space.
1.4. Unbounded Toeplitz operators on $H^{p}$
Sarason [21] demonstrated how $\mathcal{H}(b)$ spaces appear naturally as the domains of some unbounded Toeplitz operators. Let $\varphi$ be holomorphic in $\mathbb{D}$ and $T_{\varphi}$ the operator of multiplication by $\varphi$ on the domain
(1.2) $\mathrm{dom}(T_{\varphi})=\{f\in H^{2}:\varphi f\in H^{2}\}.$
Then $T_{\varphi}$ is a closed operator, and $\mathrm{dom}(T_{\varphi})$ is dense in $H^{2}$ if and only if $\varphi\in N^{+}$ (see [21, Lemma 5.2]). In this case its adjoint $T_{\varphi}^{*}$ is also densely defined and closed. In fact the domain of $T_{\varphi}^{*}$ is a de Branges-Rovnyak space.
Theorem 5.
(See [21, Prop. 5.4]) Let $\varphi$ be a nonzero function in $N^{+}$ with $\varphi=b/a$, where $(b,a)$ is the associated pair. Then $\mathrm{dom}(T_{\varphi}^{*})=\mathcal{H}(b)$.
Choosing the symbol $\varphi(z)=\frac{1}{\zeta-z}$ in Theorem 5 in conjunction with Theorem 4 gives $\mathrm{dom}(T_{\varphi}^{*})=\mathcal{D}_{\zeta}$ which played a key role in the proof of Theorem 2 (see [19]). Our goal here is to extend these ideas to $H^{p}$ spaces for all $p>1$. Let $\varphi\in N^{+}$ and define the analytic Toeplitz operator on $H^{p}$ with symbol $\varphi$ by
$T_{\varphi}f=\varphi f,\ \ \mathrm{where}\quad f\in\mathrm{dom}_{p}\,(T_{\varphi}):=\{f\in H^{p}\,:\,\phi f\in H^{p}\}.$
These $T_{\varphi}$ are bounded on $H^{p}$ precisely when $\varphi\in H^{\infty}$ (see the survey article [23]). For $\varphi=\frac{b}{a}\in N^{+}$ with $a,b\in H^{\infty}$ and $a$ outer as usual, these $T_{\varphi}$ are densely defined on $H^{p}$ for $p>1$. Indeed, $\mathrm{dom}_{p}(T_{\varphi})$ contains the dense subspace $aH^{p}$ since $a$ is outer and $T_{\varphi}(aH^{p})=bH^{p}\subset H^{p}$. It follows then that the adjoint $T_{\varphi}^{*}$ is well-defined on the dual $(H^{p})^{*}=H^{q}$ where $\frac{1}{p}+\frac{1}{q}=1$. The domain of $T_{\varphi}^{*}$ is then defined by
$\mathrm{dom}_{q}(T_{\varphi}^{*}):=\{g\in H^{q}:\exists\ h\in H^{q}\ s.t\ \langle f,h\rangle=\langle\varphi f,g\rangle\ \ \forall\ f\in\mathrm{dom}_{p}\,(T_{\varphi})\}$
[…]
Lemma 6.
Given $\varphi=\frac{b}{a}\in N^{+}$ as described above, a function $g\in\mathrm{dom}_{q}(T_{\varphi}^{*})$ if and only if there exists an $h\in H^{q}$ such that $T_{b}^{*}g=T_{a}^{*}h$.
Proof.
Suppose $g\in\mathrm{dom}_{q}(T_{\varphi}^{*})$. Then $\langle f,h\rangle=\langle\varphi f,g\rangle$ for some $h\in H^{q}$ and for all $f\in aH^{p}\subset\mathrm{dom}_{p}\,(T_{\varphi})$. Writing $f=a\tilde{f}$ for $\tilde{f}\in H^{p}$, we get
$\langle f,h\rangle=\langle\varphi f,g\rangle\ \Leftrightarrow\ \langle f,h\rangle=\langle bf/a,g\rangle\ \Leftrightarrow\ \langle a\tilde{f},h\rangle=\langle b\tilde{f},g\rangle\ \forall\ \tilde{f}\in H^{p}$
which is equivalent to $T^{*}_{a}h=T^{*}_{b}g$. This argument works in both directions because $a$ is an $H^{\infty}$ outer function and hence $aH^{p}$ is dense in $H^{p}$. ∎
We can now extend the identity $\mathrm{dom}(T_{\varphi}^{*})=\mathcal{D}_{\zeta}$ from $H^{2}$ to all $H^{p}$ with $p>1$.
Proposition 7.
Let $\varphi(z)=\frac{1}{\zeta-z}$ for $\zeta\in\mathbb{T}$ and $T_{\varphi}$ the densely defined Toeplitz operator on $H^{p}$ for any $p>1$. We then have
$\mathrm{dom}_{q}(T_{\varphi}^{*})=\mathcal{D}_{\zeta}^{q},\ \ \mathrm{where}\ \ \frac{1}{p}+\frac{1}{q}=1.$
Proof.
Choosing $\varphi=\frac{b}{a}$ with $b(z)=1$ and $a(z)=\zeta-z$ (which is outer) in Lemma 6, we get $g\in\mathrm{dom}_{q}(T_{\varphi}^{*})$ if and only if $g=T_{\zeta-z}^{*}h=(\overline{\zeta}I-S^{*})h$ for some $h\in H^{q}$. Therefore, it suffices to verify that $(\overline{\zeta}I-S^{*})H^{q}=\mathcal{D}_{\zeta}^{q}$. For any $h\in H^{q}$, we have
$(\overline{\zeta}I-S^{*})h=(\zeta I-S)(-\overline{\zeta}S^{*}h)+\overline{\zeta}h(0)\in\mathcal{D}_{\zeta}^{q}:=(S-\zeta I)H^{q}+\mathbb{C}$
and therefore $(\overline{\zeta}I-S^{*})H^{q}\subset\mathcal{D}_{\zeta}^{q}$. Conversely, if $h\in H^{q}$ and $c\in\mathbb{C}$ then
$(\zeta I-S)h+c=(\overline{\zeta}I-S^{*})\zeta(c-Sh)\in(\overline{\zeta}I-S^{*})H^{q}$
and hence $\mathcal{D}_{\zeta}^{q}\subset(\overline{\zeta}I-S^{*})H^{q}$ which concludes the proof. ∎
1.5. Cauchy duality
Let $X$ be a complete metrizable linear subspace of $\mathrm{Hol}(\mathbb{D})$. Inspired by terminology used by Malman and Seco [16], we call $X^{*}$ the Cauchy dual of $X$ if any continuous linear functional on $X$ can be represented by the Cauchy pairing
$\langle f,g\rangle:=\lim_{r\to 1^{-}}\int_{\mathbb{T}}f(r\zeta)\overline{g(r\zeta)}\,\mathrm{d}m(\zeta)\,,\quad f\in X,\ g\in X^{*}\,.$
If $H^{2}\subset X$ then $X^{*}\subset H^{2}$ and vice-versa. Hence when both $f$ and $g$ are in $H^{2}$, the pairing above reduces to the standard inner product in $H^{2}$. Some examples of Cauchy duals for our context are listed below (see [10],[15],[24]).
  1. (1)
    $H^{p}$ and $H^{q}$ for $p>1$ and $11/p+1/q=$ 1,
  2. (2)
    $H^{1}$ and $\mathrm{BMOA}$ (analytic functions with bounded mean oscillation on $\mathbb{T}$),
  3. (3)
    $H^{p}$ for $11/2<p<$ 1 and $\Lambda_{\alpha}$ (the Lipschitz class of Hol$(\mathbb{D})$-functions with $\alpha$-Hölder continuous extension to $\mathbb{T}$, where $\alpha=1/p-1$).
  4. (4)
    $N^{+}$ and the Gevrey class $\mathcal{G}$ (Hol$(\mathbb{D})$-functions whose Taylor coefficients satisfy $a_{n}=O(e^{-c\sqrt{n}})$ for some constant $c>0$).
A deep result of Davis and McCarthy [9] shows that the class $\mathcal{G}$ coincides with the universal multipliers for all non-extreme $\mathcal{H}(b)$ spaces. In particular $\mathcal{G}\subset\mathcal{H}(b)$ for all non-extreme $b$. The concept of Cauchy duality leads to an equivalence between orthogonality and density questions involving $\mathcal{N}$ which is explored in Section 2.
1.6. Minimality and biorthogonality
Let $\mathcal{H}$ be a Hilbert space. Two sequences $(e_{n})_{n\in\mathbb{N}}$ and $(f_{n})_{n\in\mathbb{N}}$ in $\mathcal{H}$ are said to be biorthogonal to each other if
$\left\langle e_{n},f_{m}\right\rangle=\delta_{nm}\ \ \ \forall\ n,m\in\mathbb{N}$
where $\delta_{nm}$ is the Kronecker delta. The sequence ($e_{n})_{n\in\mathbb{N}}$ is called minimal if $e_{n}\notin\overline{\mathrm{span}}(e_{k})_{k\neq n}$ for all $n\in\mathbb{N}$. The notions of biorthogonality, minimality and completeness are all related via the following well-known result.
Proposition 8.
(see [7, Lemma 3.3.1]) Let $(e_{n})_{n\in\mathbb{N}}$ be a sequence in $\mathcal{H}$. Then,
  • (i)
    $(e_{n})_{n\in\mathbb{N}}$ has a biorthogonal sequence if and only if $(e_{n})_{n\in\mathbb{N}}$ is minimal.
  • (ii)
    If $(e_{n})_{n\in\mathbb{N}}$ has a biorthogonal sequence, then $(e_{n})_{n\in\mathbb{N}}$ is complete in $\mathcal{H}$ if and only if its biorthogonal sequence is unique.
In Section 3 we shall prove that the sequence $(u_{k})_{k\geq 2}$ defined by
$u_{k}(z)=\sum_{d|k}\frac{\mu(k/d)}{k/d}(z^{d-1}-z^{d})$
forms a complete biorthogonal sequence for $(h_{k})_{k\geq 2}$ in $H^{2}$, where $\mu$ denotes the Möbius function defined on $\mathbb{N}$ by $\mu(k)=(-1)^{s}$ if $k$ is the product of $s$ distinct primes, and $\mu(k)=0$ otherwise.
1.7. The zeta kernels
Let $X\subset\mathrm{Hol}(\mathbb{D})$ be a topological vector space where the monomials $(z^{k})_{k\in\mathbb{N}}$ form a Schauder basis. It was shown in [14] that for each $s\in\mathbb{C}\setminus\{0\}$ a linear functional $\Lambda^{(s)}$ can be defined on $X$ by assigning
$\Lambda^{(s)}(1)=-\frac{1}{s},\quad\Lambda^{(s)}(z^{k})=-\frac{1}{s}\left((k+1)^{1-s}-k^{1-s}\right)\quad(k\geq 1).$
In particular $\Lambda^{(s)}$ is bounded on $H^{p}$ for $1<p\leq 2$ if $\Re s>1/p$ and on $H^{1}$ if $\Re s\geq 1$ (see [14, Prop. 4.7]). So there exist functions $\kappa_{s}\in H^{q}$ with $11/p+1/q=$ 1 such that $\Lambda^{(s)}(f)=\langle f,\kappa_{s}\rangle$. The function $\kappa_{s}$ will be called the zeta kernel at $s$ and
(1.3) $\kappa_{s}(z)=\sum_{k=0}^{\infty}\phi_{k}(\bar{s})z^{k}\ \ \mathrm{where}\ \ \phi_{k}(s):=\Lambda^{(s)}(z^{k}).$
The name comes from their relation to $h_{k}$ and $\zeta$ via the important identity
(1.4) $\Lambda^{(s)}(h_{k})=\langle h_{k},\kappa_{s}\rangle=-\frac{\zeta(s)}{s}(k^{1-s}-1)\ \ \forall\ \ \Re{s}>1/2,\ k\geq 2.$
The identity (1.4) appears in [14] but we provide an alternate proof in the appendix for the sake of completeness. It is important to mention that the definition of $h_{k}$ in [14] has an additional factor of $1/k$ which has been adjusted in (1.4) accordingly. The zeta kernels are used in Chapters 3 and play a key role in Chapter 4.
2. The orthogonality question
The objective of this section is to develop a framework for proving when
(2.1) $\mathcal{N}^{\perp}\cap L=\{0\}$
for topological vector spaces $L\subset H^{2}$. Since $\mathcal{N}^{\perp}=\{0\}$ is equivalent to the RH by Theorem 1, one may also ask if solutions to (2.1) can lead to new zero-free half-planes for the $\zeta$-function. We start by showing that Cauchy duality serves as a bridge between this orthogonality question and completeness questions.
Proposition 9.
Let $X$ be a topological linear space with $H^{2}\subset X\subset\mathrm{Hol}(\mathbb{D})$, where the inclusions are continuous. If $\mathcal{N}$ is dense in $X$, then
$\mathcal{N}^{\perp}\cap X^{*}=\{0\}$
where $X^{*}\subset H^{2}$ is the Cauchy dual of $X$. The converse holds when $X$ is Fréchet.
Proof.
First note that since both $\mathcal{N}^{\perp}$ and $X^{*}$ are subspaces of $H^{2}$, their intersection above makes sense. Let $\langle f,g\rangle$ denote the Cauchy pairing for $f\in X$ and $g\in X^{*}$ and recall that this pairing becomes the usual $H^{2}$-inner product $\langle f,g\rangle_{H^{2}}$ when $f,g\in H^{2}$ (see Subsection 1.5). Therefore if $\mathcal{N}$ is dense in $X$, then
$g\in\mathcal{N}^{\perp}\cap X^{*}\implies\langle f,g\rangle=\langle f,g\rangle_{H^{2}}=0\ \ \mathrm{for}\ \mathrm{all}\ \ f\in\mathcal{N}$
which implies that $g$ must be identically zero in $X^{*}$. Conversely if $X$ is additionally a Fréchet space, then we have access to the Hahn-Banach Theorem. Indeed if $\mathcal{N}$ is not dense in $X$, then there exists a non-zero $g\in X^{*}$ such that $\langle f,g\rangle=0\ \forall\,f\in\mathcal{N}$. This implies that $g$ is $H^{2}$-orthogonal to $\mathcal{N}$ and hence $g\in\mathcal{N}^{\perp}\cap X^{*}\neq\{0\}$. ∎
Since $\mathcal{N}$ is dense in $H^{p}$ for $10<p<$ 1 (see [14, Cor. 4.6]), it is also dense in $N^{+}$ since $H^{p}\subset N^{+}$ for all $p>0$. Therefore it follows by Proposition 9 that $\mathcal{N}^{\perp}\cap L=\{0\}$ if $L$ is the Lipschitz class $\Lambda_{\alpha}$ ($11/2<p<$ 1 and $\alpha=1/p-1$) or the Gevrey class $\mathcal{G}$ (see Subsection 1.5). However to obtain new zero-free half-planes for $\zeta$, we need $L\subset H^{2}$ to be large enough to contain some $H^{q}$ space for $q\geq 2$.
Corollary 10.
If $\mathcal{N}^{\perp}\cap H^{q}=\{0\}$ for some $q\geq 2$, then $\zeta(s)\neq 0$ for $\Re{s}>1/p$, where $11/p+1/q=$ 1. If $\mathcal{N}^{\perp}\cap\mathrm{BMOA}=\{0\}$, then $\zeta(s)\neq 0$ for $\Re{s}\geq 1$.
Proof.
Notice that $H^{q}$ is the Cauchy dual of $H^{p}$ and $q\geq 2$ implies that $1<p\leq 2$. In this range the $H^{p}$ are Banach spaces, and in particular Fréchet spaces. Similarly the BMOA space is the Cauchy dual of $H^{1}$. Therefore the result follows by the converse in Proposition 9 and by Theorem 1. ∎
The next result relates Toeplitz operators on $H^{p}$ and the orthogonality question. Recall that if $\varphi\in N^{+}$ is a unit, then $1/\varphi\in N^{+}$ and hence both $T_{\varphi}$ and its inverse $T_{1/\varphi}$ are densely defined Toeplitz operators on $H^{p}$ for $p>1$ (see Subsection 1.4).
Proposition 11.
Let $\varphi$ be a unit in $N^{+}$ with $T_{\varphi}$ the Toeplitz operator on $H^{p}$ for some $p>1$. If $T_{\varphi}\mathcal{N}=\varphi\mathcal{N}$ is dense in $H^{p}$, then
$\mathcal{N}^{\perp}\cap\mathrm{dom}_{q}(T_{1/\varphi}^{*})=\{0\},\ \ \ where\ \frac{1}{p}+\frac{1}{q}=1.$
The Hilbertian case $p=2$ gives $\mathcal{N}^{\perp}\cap\mathcal{H}(b)=\{0\}$ where $(b,a)$ is the Pythagorean pair associated with $1/\varphi\in N^{+}$.
Proof.
Let $g\in\mathcal{N}^{\perp}\cap\mathrm{dom}_{q}(T_{1/\varphi}^{*})$. Since both $g$ and $h_{k}$ belong to $H^{2}$ for all $k\geq 2$, the Cauchy duality $\langle h_{k},g\rangle$ coincides with the $H^{2}$-inner product $\langle h_{k},g\rangle_{2}$. Hence
$\langle\varphi h_{k},T^{*}_{1/\varphi}g\rangle=\langle h_{k},g\rangle=\langle h_{k},g\rangle_{2}=0\ \ \forall\ k\in\mathbb{N}$
which implies that $T^{*}_{1/\varphi}g=0$ in $H^{q}$ by the density of $\varphi\mathcal{N}$ in $H^{p}$. Since $1/\varphi=b/a$ is a unit in $N^{+}$ (as the inverse of $\varphi$), both $a$ and $b$ are $H^{\infty}$ outer functions in $H^{p}$. So $T_{1/\varphi}(aH^{p})=bH^{p}$ shows that $T_{1/\varphi}$ has dense range in $H^{p}$ since $b$ is outer. So $T^{*}_{1/\varphi}$ is injective and therefore $g=0$ in $H^{q}$. This concludes the general case. The Hilbertian case $p=2$ now follows by Theorem 5. ∎
We shall derive two non-trivial applications of this result. The first one extends Theorem 2 to all local Dirichlet spaces $\mathcal{D}_{1}^{p}$ with $p>1$. We note that the classical $\mathcal{D}_{\delta_{1}}$ is just $\mathcal{D}_{1}^{2}$ which is strictly smaller than $\mathcal{D}_{1}^{p}$ for $p\in(1,2)$ (see Subsection 1.2). We will need with the following approximation result from [14].
Lemma 12.
Let $\mu$ the Möbius function. Then
$\sum_{k=2}^{n}\frac{\mu(k)}{k}(I-S)h_{k}\to 1-z$
in the $H^{p}$ norm for all $p>0$.
We observe that $I-S=T_{\varphi}$ where $\varphi(z)=1-z$ is an $H^{\infty}$ outer function since $\mathbb{C}[z]\varphi$ is dense in $H^{2}$. In particular $\varphi$ is a unit in $N^{+}$. Define operators on $H^{p}$ by
(2.2) $(W_{n})f(z)=(1+z+\dots+z^{n-1})f(z^{n})=\frac{1-z^{n}}{1-z}f(z^{n})$
and $(T_{n})f(z)=f(z^{n})$ for $n\geq 1$ and $f\in H^{p}$. The multiplicative semigroup of operators $(W_{n})_{n\geq 1}$ was introduced in [19] and is the main object of study in [17]. They are bounded on $H^{p}$ for $p>1$ (see [14, Cor. 4.6]). We shall need the identities
$W_{n}\mathcal{N}\subset\mathcal{N}\ \ \mathrm{and}\ \ T_{n}(I-S)=(I-S)W_{n}$
for $k,n\geq 1$ which appear in [19, p. 249]. We are ready for the first application.
Theorem 13.
We have $\mathcal{N}^{\perp}\cap\mathcal{D}_{1}^{p}=\{0\}$ for all $p>1$.
Proof.
Let $\varphi(z)=1-z$. By Propositions 7 and 11 we only need to prove that $\varphi\mathcal{N}$ is dense in $H^{p}$ for $p>1$. First note that Lemma 12 implies that $\varphi$ belongs to the $H^{p}$-closure of $\varphi\mathcal{N}=(I-S)\mathcal{N}$. This in turn implies that $T_{n}\varphi$ belongs to the $H^{p}$-closure of $T_{n}(\varphi\mathcal{N})=\varphi W_{n}\mathcal{N}\subset\varphi\mathcal{N}$ for all $n\geq 1$. So in particular $\mathrm{span}(T_{n}\varphi)_{n\geq 1}\subset\ \mathrm{clos}_{H^{p}}(\varphi\mathcal{N})$. Now $\mathrm{span}(T_{n}\varphi)_{n\geq 1}=\mathrm{span}(1-z^{n})_{n\geq 1}=\mathbb{C}[z]\varphi$ which is dense in $H^{p}$ for all $p>1$ because $\varphi$ is an $H^{\infty}$ outer function. This proves that $\mathrm{clos}_{H^{p}}(\varphi\mathcal{N})=H^{p}$ and concludes the proof. ∎
Our second application of Proposition 11 utilizes recent discoveries in $\mathcal{H}(b)$-space theory to obtain zero-free half-planes for $\zeta$. In view of Corollary 10, we would like to know when the $H^{p}$ and BMOA spaces are contained in some $\mathcal{H}(b)$ for $\varphi=b/a\in N^{+}$. Fortuitously for us, these problems were completely solved recently in a preprint by Malman and Seco [16]. They show that $H^{\tilde{p}}\subset\mathcal{H}(b)$ for $\tilde{p}\in(2,\infty)$ if and only if $\varphi\in H^{p}$ where $p=\frac{2\tilde{p}}{\tilde{p}-2}\in(2,\infty)$, and also that $H^{\infty}\subset\mathrm{BMOA}\subset\mathcal{H}(b)$ if and only if $\varphi\in H^{2}$. By definition we always have $\mathcal{H}(b)\subset H^{2}$, and $\mathcal{H}(b)=H^{2}$ precisely when $\varphi\in H^{\infty}$. Therefore it makes sense to allow the values $p=2$ and $p=\infty$.
Theorem 14.
Suppose $\varphi$ is a unit in $N^{+}$ such that $1/\varphi\in H^{p}$ for some $p\in(2,\infty)$. If $\varphi\mathcal{N}$ is dense in $H^{2}$, then $\zeta(s)\neq 0$ for $\Re{s}>\frac{1}{2}+\frac{1}{p}$. The case $p=2$ gives the Prime Number Theorem ($\Re{s}\geq 1$) and $p=\infty$ gives the RH ($\Re{s}>1/2$).
Proof.
By the results of Malman and Seco [16] mentioned above, $1/\varphi$ belongs to $H^{2}$ or to $H^{\infty}$ precisely when $\mathcal{H}(b)$ contains BMOA or $\mathcal{H}(b)=H^{2}$ respectively, where $1/\varphi=b/a$ and $(b,a)$ the associated Pythagorean pair. Hence the cases $p=2,\infty$ follow by Corollary 10 and Proposition 11. For the case when $1/\varphi\in H^{p}$ for $p\in(2,\infty)$, we have $H^{\tilde{p}}\subset\mathcal{H}(b)$ where $p=\frac{2\tilde{p}}{\tilde{p}-2}$ (again by Malman and Seco) or equivalently $\tilde{p}=\frac{2p}{p-2}$. If $11/\tilde{p}+1/\tilde{q}=$ 1, then we see that $\tilde{q}=\frac{2p}{(p+2)}$ and hence $1/\tilde{q}=1/2+1/p$. The result again follows by Corollary 10 and Proposition 11. ∎
An important distinction between Theorem 1 and Theorem 14 is that in the former one must solve density problems in $H^{p}$ spaces that are non-Hilbertian, while in the latter the density problems are always in $H^{2}$. The following simple examples of $\varphi$ satisfy the hypothesis of Theorem 14. Let $\varphi(z)=(1-z)^{\alpha}$ for some $20<\alpha<1/$ 2. Then $\varphi$ is an $H^{\infty}$ outer function and hence a unit in $N^{+}$ with the property that $1/\varphi\in H^{p}$ for some $2<p<1/\alpha$. It follows that the density of $\varphi\mathcal{N}$ in $H^{2}$ would give a new zero-free half-plane for $\zeta$.
3. The biorthogonality question
Define the sequence of polynomials $\{u_{k}:k\geq 2\}$ by
(3.1) $u_{k}(z)=\sum_{d|k}\frac{\mu(k/d)}{k/d}(z^{d-1}-z^{d}),$
where $\mu$ denotes the Möbius function and $d|k$ denotes $d$ divides $k$. The main goal of this section is to prove the following theorem.
Theorem 15.
$\{u_{k}:k\geq 2\}$ is complete and biorthogonal to $\{h_{k}:k\geq 2\}$ in $H^{2}$.
Balazard [5] noted that with the additional vector $u_{1}(z)=1-z$ the sequence $\{u_{k}:k\geq 1\}$ is complete. However it is no longer minimal following Theorem 15. We first make a key observation. Note that $u_{k}=(I-S)v_{k}$, where
(3.2) $v_{k}(z)=\sum_{d|k}\frac{\mu(k/d)}{k/d}z^{d-1}.$
It follows that $\left\langle h_{k},\,u_{j}\right\rangle=\left\langle(I-S^{*})h_{k},v_{j}\right\rangle.$ Hence to show that $\{u_{k}:k\geq 2\}$ and $\{h_{k}:k\geq 2\}$ are biorthogonal, it is suffices to show that
$\left\langle(I-S^{*})h_{k},v_{j}\right\rangle=\delta_{kj}.$
The proof of Theorem 15 will be divided into four steps.
[…]
Lemma 16.
We have
$(I-S^{*})h_{k}(z)=\sum_{n=0}^{\infty}B_{k}(n+1)z^{n}$
for all $k\geq 2$ where
$B_{k}(n)=\begin{cases}\frac{k}{n}-\frac{1}{n},\quad k|n\\ -\frac{1}{n},\quad k\!\!\not|n\end{cases}.$
Proof.
Note that if $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$, then
$S^{*}f(z)=\sum_{n=0}^{\infty}a_{n+1}z^{n}.$
Let $c_{n}(k)$ be the Fourier coefficients of $h_{k}$, i.e.,
$h_{k}(z)=\sum_{n=0}^{\infty}c_{n}(k)z^{n}.$
Then, the $n$-th Fourier coefficient of $(I-S^{*})h_{k}$ is $c_{n}(k)-c_{n+1}(k)$. The coefficients $c_{n}(k)$ are calculated in [19, p. 249]:
(3.3) $c_{n}(k)=H(n)-H\left(\frac{n}{k}\right)-\log k,$
where $H(x):=\sum_{n\leq x}\frac{1}{n}$ for $x>0$ and $H(0)=0$. It follows from (3.3) that
$\begin{split}c_{n-1}(k)-c_{n}(k)&=H(n-1)-H\left(\frac{n-1}{k}\right)-\log k-\left(H(n)-H\left(\frac{n}{k}\right)-\log k\right)\\ &=-\frac{1}{n}+\sum_{\frac{n-1}{k}<m\leq\frac{n}{k}}\frac{1}{m}.\end{split}$
Note that if there is some $m\in\mathbb{N}$ such that $\frac{n-1}{k}<m\leq\frac{n}{k}$, then $mk\leq n<mk+1$, so that $n=mk$. Therefore, the sum above is non-zero if and only if $k|n$. Then,
(3.4) $B_{k}(n)=c_{n-1}(k)-c_{n}(k)=\begin{cases}-\frac{1}{n},\quad k\!\!\not|n\\ \frac{k}{n}-\frac{1}{n},\quad k|n\end{cases}.$
[…]
Theorem 17.
$\{u_{k}:k\geq 2\}$ is biorthogonal to $\{h_{k}:k\geq 2\}$.
Proof.
[…]
$\sum_{d|j}B_{k}(d)\frac{\mu(j/d)}{j/d}=\left\langle(I-S^{*})h_{k},v_{j}\right\rangle=\delta_{kj},\qquad\forall k,j\geq 2.$
There are two cases:
  • (i)
    $k\!\!\not|j$. Then, $k\!\!\not|d$ for every $d|j$, therefore
    $\sum_{d|j}B_{k}(d)\frac{\mu(j/d)}{j/d}=\sum_{d|j}-\frac{1}{d}\frac{\mu(j/d)}{j/d}=-\frac{1}{j}\sum_{d|j}\mu(j/d)=-\frac{1}{j}\left\lfloor\frac{1}{j}\right\rfloor=0,$
    since $j\geq 2$ and by the basic relation $\sum_{d|k}\mu(d)=\lfloor 1/k\rfloor$.
  • (ii)
    $k|j$. Let $q=\frac{j}{k}$. Then
    $\sum_{d|j}B_{k}(d)\frac{\mu(j/d)}{j/d}=\sum_{\begin{subarray}{c}d|j\\ k\nmid d\end{subarray}}-\frac{1}{d}\frac{\mu(j/d)}{j/d}+\sum_{\begin{subarray}{c}d|j\\ k|d\end{subarray}}\left(\frac{k}{d}-\frac{1}{d}\right)\frac{\mu(j/d)}{j/d}.\$
The last sum is summing over those $d$ that satisfy $k|d|j$. However, $k|d\iff d=mk$ for some $m\in\mathbb{N}$. Since $j=qk$, it follows that $d|j\iff m|q$. Hence, the last sum can be written as
$\sum_{\begin{subarray}{c}d|j\\ k|d\end{subarray}}\left(\frac{k}{d}-\frac{1}{d}\right)\frac{\mu(j/d)}{j/d}=\sum_{\begin{subarray}{c}d|j\\ k|d\end{subarray}}-\frac{1}{d}\frac{\mu(j/d)}{j/d}+\sum_{m|q}\frac{1}{m}\frac{\mu(q/m)}{q/m}.$
Therefore,
$\sum_{d|j}B_{k}(d)\frac{\mu(j/d)}{j/d}$ $=\sum_{d|j}-\frac{1}{d}\frac{\mu(j/d)}{j/d}+\sum_{m|q}\frac{1}{m}\frac{\mu(q/m)}{q/m}$
$=-\frac{1}{j}\sum_{d|j}\mu(j/d)+\frac{1}{q}\sum_{m|q}\mu(q/m)=-\frac{1}{j}\left\lfloor\frac{1}{j}\right\rfloor+\frac{1}{q}\left\lfloor\frac{1}{q}\right\rfloor.$
Since $j\geq 2$, the first term is always 0. On the other hand, the second term equals 1 if $q=1$ and equals 00 otherwise. Finally note that $q=1\iff k=j$, and hence that $\left\langle h_{k},u_{j}\right\rangle=\delta_{kj}$ for all $k,j\geq 2$. ∎
[…]
Lemma 18.
A sequence $\{f_{k}:k\geq 2\}\subset H^{2}$ is biorthogonal to $\{v_{k}:k\geq 2\}$ if and only if there exists a sequence $(c_{k})_{k\geq 2}\in\mathbb{C}^{\mathbb{N}}$ such that
$f_{k}(z)=\sum_{n=0}^{\infty}A_{k}(n+1)\ z^{n},\quad\forall k\geq 2,$
where the sequence $(A_{k}(n))_{n\geq 1}$ for each $k\geq 2$ is defined by
$A_{k}(n)=\begin{cases}\frac{c_{k}}{n}+\frac{k}{n},\quad k|n\\ \frac{c_{k}}{n},\ \ \ \ \ \quad k\!\!\not|n\end{cases}.$
Proof.
Let $\{f_{k}:k\geq 2\}\subset H^{2}$ be a sequence biorthogonal to $\{v_{k}:k\geq 2\}$ and $A_{k}:\mathbb{N}\to\mathbb{C}$ be the arithmetical functions that satisfy
$f_{k}(z)=\sum_{n=0}^{\infty}A_{k}(n+1)z^{n},\quad\forall k\geq 2.$
Since the coefficients of $v_{j}$ are real, the biorthogonality condition becomes
(3.5) $\sum_{d|j}A_{k}(d)\frac{\mu(j/d)}{j/d}=\left\langle f_{k},v_{j}\right\rangle=\delta_{kj},\quad\forall k,j\geq 2.$
Let $I_{k},\nu:\mathbb{N}\to\mathbb{C}$ be arithmetic functions defined by $I_{k}(n)=\delta_{kn}$ and $\nu(n)=\frac{\mu(n)}{n}$. Then (3.5) is equivalent to
(3.6) $\forall\ k\geq 2,\,\exists\ c_{k}\in\mathbb{C}\ \text{ such that}\quad A_{k}*\nu=c_{k}I_{1}+I_{k},$
where * denotes the Dirichlet product (see [1, Section 2.6]). Indeed, (3.5) doesn’t impose any restriction on $A_{k}*\nu(1)$, since it only need to hold for $j\geq 2$, hence $c_{k}=A_{k}*\nu(1)$ is free, so (3.5) and (3.6) are indeed equivalent. Notice that
$\sum_{d|k}\frac{\mu(k/d)}{k/d}\frac{1}{d}=\frac{1}{k}\left\lfloor\frac{1}{k}\right\rfloor=I_{1}(k),$
i.e., $\nu^{-1}(n)=\frac{1}{n}$, since $I_{1}$ is the unity with respect to *. Moreover
$I_{k}*\nu^{-1}(n)=\sum_{d|n}\delta_{kd}\frac{1}{n/d}=\begin{cases}\frac{k}{n},\quad k|n\\ 0,\quad k\!\!\not|n\end{cases}.$
Therefore (3.6) is equivalent to the statement that
$\forall\ k\geq 2,\,\exists\ c_{k}\in\mathbb{C}\text{ such that }A_{k}(n)$ $=c_{k}\nu^{-1}(n)+I_{k}*\nu^{-1}(n)$
=$=\begin{cases}\frac{c_{k}}{n}+\frac{k}{n},\quad k|n\\ \frac{c_{k}}{n},\ \ \ \ \ \quad k\!\!\not|n\end{cases}.$
Hence the biorthogonality condition (3.5) is equivalent to the condition above as desired. Finally $f_{k}\in H^{2}$ since its coefficient sequence $A_{k}$ clearly belongs to $\ell^{2}$. ∎
[…]
$\left\langle\phi_{k},\,u_{j}\right\rangle=\left\langle(I-S^{*})\phi_{k},v_{j}\right\rangle$
for any sequence $\{\phi_{k}:k\geq 2\}$ in $H^{2}$. This implies that $I-S^{*}$ maps sequences biorthogonal to $\{u_{k}:k\geq 2\}$ onto sequences biorthogonal to $\{v_{k}:k\geq 2\}$ in the image of $I-S^{*}$. This correspondece is one-to-one since $I-S^{*}$ is injective on $H^{2}$. Therefore it is enough to prove that $((I-S^{*})h_{k})_{k\geq 2}$ is the unique sequence in the image of $I-S^{*}$ that is biorthogonal to $\{v_{k}:k\geq 2\}$.
Lemma 19.
A sequence $\{f_{k}:k\geq 2\}\subset(I-S^{*})H^{2}$ is biorthogonal to $\{v_{k}:k\geq 2\}$ if and only if
(3.7) $f_{k}(z)=\sum_{n=0}^{\infty}B_{k}(n+1)z^{n}=(I-S^{*})h_{k},$
where $B_{k}$ are the sequences defined in Lemma 16.
Proof.
Let $\{f_{k}:k\geq 2\}\subset(I-S^{*})H^{2}$ be a sequence biorthogonal to $\{v_{k}:k\geq 2\}$ and let $\varphi_{k}\in H^{2}$ such that $f_{k}=(I-S^{*})\varphi_{k}$. If $(b_{k}(n))_{n\geq 0}$ are the Maclaurin coefficients of $\varphi_{k}$, then
$f_{k}(z)=\sum_{n=0}^{\infty}(b_{k}(n)-b_{k}(n+1))z^{n}.$
It then follows by Lemma 18 that for each $k\geq 2$, there exists a $c_{k}\in\mathbb{C}$ such that
$b_{k}(n-1)-b_{k}(n)=A_{k}(n)=\begin{cases}\frac{k}{n}+\frac{c_{k}}{n},\quad k|n\\ \frac{c_{k}}{n},\quad k\!\!\not|n\end{cases},\quad\forall n\geq 1.$
By induction, we obtain
$b_{k}(n)$ $=b_{k}(0)-\sum_{j=1}^{n}A_{k}(j)=b_{k}(0)-\sum_{j\leq n}\frac{c_{k}}{j}-\sum_{\begin{subarray}{c}j\leq n\\ k|j\end{subarray}}\frac{k}{j}$
$=b_{k}(0)-c_{k}\sum_{j\leq n}\frac{1}{j}-\sum_{m\leq n/k}\frac{1}{m}=b_{k}(0)-c_{k}H(n)-H\left(\frac{n}{k}\right),$
where $H$ is the same function used in (3.3). Since $\varphi_{k}\in H^{2}$, we get $(b_{k}(n))_{n}\in\ell^{2}$ and hence $\lim_{n\to\infty}b_{k}(n)=0$. So $(c_{k}H(n)+H(n/k))_{n}$ converges. Using Euler summation, one gets (see [19])
(3.8) $H(x)=\log x+\gamma+O\left(\frac{1}{x}\right),$
where $\gamma$ is Euler-Mascheroni constant. Therefore,
$c_{k}H(n)+H\left(\frac{n}{k}\right)$ $=c_{k}\log n+c_{k}\gamma+\log n-\log k+\gamma+O\left(\frac{k}{n}\right)$
$=(c_{k}+1)\log n+(c_{k}+1)\gamma-\log k+O\left(\frac{k}{n}\right),$
which converges as $n\to\infty$ if and only if $c_{k}=-1$. Hence $c_{k}=-1$ for all $k\geq 2$. In that case $A_{k}=B_{k}$ and we obtain (3.7). The converse is equivalent to Theorem 17 by the remarks at the start of Step $4$. ∎
As a consequence of Theorem 1 and Proposition 8 the RH holds if and only if $(u_{j})_{j\geq 2}$ is the unique sequence in $H^{2}$ that is biorthogonal to $(h_{k})_{k\geq 2}$. On the other hand the next result shows what happens if some $\zeta$-zero violates the RH.
Corollary 20.
If $\zeta(s_{0})=0$ for some $11/2<\Re{s_{0}}<$ 1, then
$\left\langle h_{k},u_{j}+\kappa_{s_{0}}\right\rangle=\delta_{kj}\ \ \forall\ \ k,j\geq 2$
where $\kappa_{s_{0}}$ is the zeta kernel at $s_{0}$. So $(u_{j}+\kappa_{s_{0}})_{j\geq 2}$ is also biorthogonal to $(h_{k})_{k\geq 2}$.
Proof.
This follows by (1.4) and Theorem 17 since $\langle h_{k},\kappa_{s_{0}}\rangle=0$ for all $k\geq 2$. ∎
4. The RH-failure conjecture
The RH-failure (RHF) conjecture states that if the RH is false, then $\zeta(s)=0$ for infinitely many $s\in\mathbb{C}$ with $11/2<\Re{s}<$ 1. Our goal is to prove the following.
Theorem 21.
The RHF conjecture implies that $\dim(\mathcal{N}^{\perp})$ is either 0 or $\infty$.
Let $\mathcal{K}:=\{\kappa_{s}:\Re{s}>1/2\}$ denote the family of zeta kernels. If $\zeta(s)=0$ for some $\Re{s}>1/2$, then $\langle h_{k},\kappa_{s}\rangle=0$ for all $k\geq 2$ by (1.4) and hence $\kappa_{s}\in\mathcal{N}^{\perp}$ So the RHF conjecture implies that $\mathcal{N}^{\perp}\cap\mathcal{K}$ is either empty (by Theorem 1) or has infinitely many elements. Therefore Theorem 21 follows if we show that $\mathcal{K}$ is linearly independent in $H^{2}$. We first show that elements of $\mathcal{K}$ are common eigenvectors for the adjoints of operators $(W_{n})_{n\geq 1}$ defined in (2.2). For $f\in H^{2}$ and $n\in\mathbb{N}$, we have
$W^{*}_{n}f(z)=\sum_{k=0}^{\infty}[\hat{f}(nk)+\hat{f}(nk+1)+\ldots+\hat{f}(nk+n-1)]z^{k}$
where $\hat{f}(n)$ denotes the $n$-th Fourier coefficient of $f$. This formula first appeared in [17]. It is possible to describe the common eigenvectors of $(W_{n}^{*})_{n\geq 1}$ completely.
Proposition 22.
A non-zero $f\in H^{2}$ is a common eigenvector for $(W_{n}^{*})_{n\geq 1}$ if and only if there exists a multiplicative sequence $(\lambda_{n})_{n\geq 1}$ with $(\lambda_{n+1}-\lambda_{n})_{n\geq 1}\in\ell^{2}$ and
(4.1) $\hat{f}(n)=(\lambda_{n+1}-\lambda_{n})\hat{f}(0)\ \ \forall\ \ n\geq 1.$
Moreover $W_{n}^{*}f=\lambda_{n}f$ for all $n\geq 1$.
By a multiplicative sequence $(\lambda_{n})_{n\geq 1}$ we mean that $\lambda_{n}\lambda_{m}=\lambda_{nm}$ and $\lambda_{1}=1$. Similarly one can see that $W_{n}W_{m}=W_{nm}$ and $W_{1}=I$ by (2.2).
Proof.
Let $W_{n}^{*}f=\lambda_{n}f$ for $n\geq 1$ and some sequence $(\lambda_{n})_{n\geq 1}$. Since $W_{1}^{*}=I$ and $W_{nm}^{*}=W_{n}^{*}W_{m}^{*}$ it follows that $(\lambda_{n})_{n\geq 1}$ is multiplicative. Furthermore
$\lambda_{n}\hat{f}(k)=\langle W_{n}^{*}f,z^{k}\rangle=\langle f,W_{n}z^{k}\rangle=\left\langle f,\sum_{j=0}^{n-1}z^{nk+j}\right\rangle=\sum_{j=0}^{n-1}\hat{f}(nk+j).$
which gives $\hat{f}(n)=\lambda_{n+1}\hat{f}(0)-\lambda_{n}\hat{f}(0)$ for all $n\geq 1$ and hence $(\lambda_{n+1}-\lambda_{n})_{n\geq 1}\in\ell^{2}$. Conversely suppose $f$ is a non-zero function satisfying (4.1) for some multiplicative $(\lambda_{n})_{n\geq 1}$ with $(\lambda_{n+1}-\lambda_{n})_{n\geq 1}\in\ell^{2}$. Normalizing by supposing $\hat{f}(0)=1$, we get
$(W_{n}^{*}f)(z)$ $=\sum_{k=0}^{\infty}\left(\sum_{j=0}^{n-1}\hat{f}(nk+j)\right)z^{k}=\sum_{j=0}^{n-1}\hat{f}(j)+\sum_{k=1}^{\infty}\left(\sum_{j=0}^{n-1}\hat{f}(nk+j)\right)z^{k}$
$=\lambda_{n}+\sum_{k=1}^{\infty}(\lambda_{nk+n}-\lambda_{nk})z^{k}=\lambda_{n}\left(1+\sum_{k=1}^{\infty}(\lambda_{k+1}-\lambda_{k})z^{k}\right)=\lambda_{n}f(z)$
for all $n\geq 2$. So $f\in H^{2}$ is a common eigenvector for $(W_{n}^{*})_{n\geq 1}$. ∎
Choosing $\lambda_{k}=k^{1-\bar{s}}$ and $\hat{f}(0)=-1/\bar{s}$ in Proposition 22 for any fixed $\Re{s}>1/2$ shows that each $\kappa_{s}\in\mathcal{K}$ is a common eigenvector for $(W_{n}^{*})_{n\geq 1}$ (see (1.3)) with
(4.2) $W_{n}^{*}\kappa_{s}=n^{1-\bar{s}}f\ \ \forall\ n\geq 1.$
We want to prove that for any finite subset $\{\kappa_{s_{1}},\ldots,\kappa_{s_{\ell}}\}\subset\mathcal{K}$ there exists some $W_{n}^{*}$ such that the corresponding eigenvalues are all distinct. This will give us the linear independence of every finite subset of $\mathcal{K}$ and hence of $\mathcal{K}$ itself. First suppose that the real parts of $s_{1},\,\dots,s_{\ell}$ are all distinct. Since $|n^{1-\bar{s}}|=n^{1-\Re{s}}$ it follows that the eigenvalues of $W_{n}^{*}$ (for all $n>1$) corresponding to $\kappa_{s_{1}},\ldots,\kappa_{s_{\ell}}$ are all distinct. If the real parts of $s_{1},\,\dots,s_{\ell}$ are not all distinct, then we need the following result.
Lemma 23.
Given distinct $a_{1},\,\dots,\,a_{n}\in\mathbb{R}$, at most finitely many primes $p$ have the property that there exists a pair $a_{i},a_{j}$ with $1\leq i<j\leq n$ such that
(4.3) $\ (a_{i}-a_{j})\log p\in 2\pi\mathbb{Z}.$
Proof.
Suppose there are infinitely many primes that satisfy (4.3). For each such prime $p$ there exists some $1\leq i<j\leq n$ and $k\in\mathbb{Z}\setminus\{0\}$ such that
$(a_{i}-a_{j})\log p=2\pi k\ \implies\ \frac{2\pi k}{\log p}=a_{i}-a_{j}.$
But since there are only finitely many numbers $a_{i}-a_{j}$ with $i<j$, and none of which equal 0, there must exist distinct primes $p,q$ and $k_{1},k_{2}\in\mathbb{Z}\setminus\{0\}$ such that
$\frac{2\pi k_{1}}{\log p}=a_{i}-a_{j}=\frac{2\pi k_{2}}{\log q}\ \implies\ k_{2}\log{p}=k_{1}\log{q}\neq 0.$
for some pair $i<j$. In particular, $p^{k_{2}}=q^{k_{1}}\neq 1$, which is a contradiction. ∎
The following result then completes the proof of Theorem 21.
Proposition 24.
The family of zeta kernels $\mathcal{K}$ is linearly independent.
Proof.
Let $\{\kappa_{s_{1}},\ldots,\kappa_{s_{\ell}}\}\subset\mathcal{K}$ be a finite subset. The case when the real parts of $s_{1},\,\dots,s_{\ell}$ are all distinct was already dealt with. Suppose some of the $s_{1},\,\dots,s_{\ell}$ have the same real parts. So $\{s_{1},\,\dots,s_{\ell}\}$ is the finite disjoint union of sets of the form $A_{r}:=\{s_{i}:\Re{s_{i}}=r,\ i=1,\ldots,\ell\}$ for $r\in\mathbb{R}$. It is enough to prove that the family $\{\kappa_{s}:s\in A_{r}\}$ is linearly independent when $A_{r}$ has more than one element. Since $s_{1},\,\dots,s_{\ell}$ are distinct complex numbers, the imaginary parts of elements in $A_{r}$, which we denote by $a_{1},\ldots,a_{n}$, must all be distinct. Applying Lemma 23 to $a_{1},\ldots,a_{n}$ shows that there exist infinitely many primes $q$ such that
[…]
For such a prime $q$, we claim that the $\{\kappa_{s}:s\in A_{r}\}$ are $W_{q}^{*}$-eigenvectors with distinct eigenvalues. To see this first note that $W_{q}^{*}\kappa_{s}=q^{1-\bar{s}}\kappa_{s}$ by (4.2) and
$q^{1-\bar{s}}=e^{(1-\bar{s})\log{q}}=e^{(1-r)\log q}e^{i\mathrm{Im}{(s)}\log q}\ \ \forall\ s\in A_{r}.$
But Im$(s)$ for $s\in A_{r}$ are precisely the real numbers $a_{1},\ldots,a_{n}$. Therefore the eigenvalues $q^{1-\bar{s}}$ for $s\in A_{r}$ are all distinct by (4.4) and hence $\{\kappa_{s}:s\in A_{r}\}$ and therefore all of $\mathcal{K}$ is linearly independent. ∎
5. Appendix
Denote by $\mathbb{C}_{\rho}$ the half-plane $\{s\in\mathbb{C}:\Re{s}>\rho\}$. In this appendix we provide an alternate proof for the fundamental relation
(5.1) $\langle h_{k},\kappa_{s}\rangle=-\frac{\zeta(s)}{s}(k^{1-s}-1)\ \ \forall\ s\in\mathbb{C}_{1/2},\ k\geq 2.$
We first prove that (5.1) holds for all $s\in\mathbb{C}_{1}$. We then prove that the function $s\longmapsto\langle h_{k},\kappa_{s}\rangle$ has an analytic continuation to $\mathbb{C}_{0}$ for each $k\geq 2$. Since the right side of (5.1) is already analytic for $s\in\mathbb{C}\setminus\{0\}$, the result then follows by analytic continuation. Recall from Subsection 1.7 that
$\kappa_{s}(z)=\sum_{n=0}^{\infty}\phi_{n}(\bar{s})z^{n}\ \ \mathrm{where}\ \ \phi_{n}(s)=-\frac{1}{s}\left((n+1)^{1-s}-n^{1-s}\right).$
Lemma 25.
The identity (5.1) holds for $s\in\mathbb{C}_{1}$.
Proof.
Let $(c_{n}(k))_{n}$ be the Fourier coefficients of $h_{k}$. Since $\overline{\phi_{n}(\overline{s})}=\phi_{n}(s)$, we have
$\langle h_{k},\kappa_{s}\rangle$ $=\sum_{n=0}^{\infty}c_{n}(k)\overline{\phi_{n}(\overline{s})}=\sum_{n=0}^{\infty}c_{n}(k)\phi_{n}(s)$
$=\lim_{N\to\infty}\left(-\frac{c_{0}(k)}{s}-\frac{1}{s}\sum_{n=1}^{N}c_{n}(k)\left((n+1)^{1-s}-n^{1-s}\right)\right)$
$=\lim_{N\to\infty}\left(-\frac{1}{s}\sum_{n=0}^{N}c_{n}(k)(n+1)^{1-s}+\frac{1}{s}\sum_{n=1}^{N}c_{n}(k)n^{1-s}\right)$
$=\lim_{N\to\infty}\left(-\frac{1}{s}\sum_{n=1}^{N}(c_{n-1}(k)-c_{n}(k))n^{1-s}-\frac{1}{s}c_{N}(k)(N+1)^{1-s}\right).$
Since $c_{n}(k)=O(k/n)$ (see [19, p. 249]), we have $c_{N}(k)(N+1)=O(1)$. Furthermore $(N+1)^{-s}\to 0$ for $\Re(s)>0$ and therefore we get
$\langle h_{k},\kappa_{s}\rangle$ $=-\frac{1}{s}\lim_{N\to\infty}\left(\sum_{n=1}^{N}(c_{n-1}(k)-c_{n}(k))n^{1-s}\right)$
=(3.4)${=}-\frac{1}{s}\lim_{N\to\infty}\left(\sum_{n=1}^{N}-\frac{1}{n}n^{1-s}+\sum_{\begin{subarray}{c}n=1\\ k|n\end{subarray}}^{N}\frac{k}{n}n^{1-s}\right)$ )
$=-\frac{1}{s}\lim_{N\to\infty}\left(\sum_{n=1}^{N}n^{-s}+\sum_{m=1}^{\lfloor\frac{N}{k}\rfloor}\frac{1}{m}(mk)^{1-s}\right)$
$=-\frac{1}{s}\lim_{N\to\infty}\left(-\sum_{n=1}^{N}n^{-s}+k^{1-s}\sum_{m=1}^{\lfloor\frac{N}{k}\rfloor}\frac{1}{m}m^{1-s}\right)$
=$\overset{(*)}{=}-\frac{1}{s}(-\zeta(s))-\frac{k^{1-s}}{s}\zeta(s)=-\frac{\zeta(s)}{s}(k^{1-s}-1),$
where in ()(*)( ∗ ) we split the limit in two and use the definition of $\zeta$ for $\Re(s)>1$. ∎
The inner product $\langle h_{k},\kappa_{s}\rangle$ defined for $s\in\mathbb{C}_{1/2}$ also makes sense for $s\in\mathbb{C}_{0}$.
Lemma 26.
The function $\Phi_{k}:\mathbb{C}_{1/2}\to\mathbb{C}$ defined by
(5.2) $\Phi_{k}(s):=\langle h_{k},\kappa_{s}\rangle=\sum_{n=0}^{\infty}c_{n}(k)\phi_{n}(s).$
has an analytic continuation to $\mathbb{C}_{0}$ for each $k\geq 2$.
Proof.
Since each $\phi_{n}$ is holomorphic in $\mathbb{C}_{0}$, it is sufficient to prove that the series in (5.2) converges uniformly in every half-plane $\mathbb{C}_{\rho}$ for $\rho>0$. Note that
[…]
for $s\in\mathbb{C}_{\rho}$ with $\rho>0$. Also $c_{n}(k)=O(k/n)$ for each $k\geq 2$, and hence we get $c_{n}(k)\phi_{n}(s)=O(n^{-1-\rho})$ for $s\in\mathbb{C}_{\rho}$. So $\Phi_{k}$ converges uniformly in $\mathbb{C}_{\rho}$ for $\rho>0$. ∎
Acknowledgement
This work was financed in part by the Coordenação de Aperfeiçoamento de Nivel Superior- Brasil (CAPES)- Finance Code 001. The fourth author is supported by grant #2023 Provost of Inclusion and Belonging, University of São Paulo (USP).